OpenStudy (anonymous):

Several teachers were meeting in the library. After 45 of them left, the room was 5/8 as full as it was initially. How many teachers were in the room at the start of the meeting?

OpenStudy (naveen):

x = number of teachers in room initially x - 45 = 5/8x x = 5/8x + 45 x - 5/8x = 45 3/8x = 45 x = 120 120 teachers

OpenStudy (anonymous):

thank you

OpenStudy (naveen):

Your welcome